A student was studying the balanced reaction: CS2+3O2=CO2+2SO2
Answer the following questions if this reaction started with 4 molecules of CS2 and 9 molecules of O2.
what chemical in this balanced reaction is the limiting reactant?
How many molecules of CO2 would be made on a complete reaction?
how many molecules of the SO2 would be made in a complete reaction?
how many molecules of CS2 would remain at the end of a complete reaction?
how many molecules of O2 would remain at the end of a complete reaction?
1
Expert's answer
2019-07-05T04:58:16-0400

1) In this reaction ratio of reagents:

"n(CS_2) : 3n(O_2)"


Respectively if  this reaction started with 4 molecules of CS2 and 9 molecules of O2 ratio (CS2) : (O2):

"4 : 9"

"1 : 2.25"

Respectively the limiting reactant is O2.

2)

"3n(O_2) : n(CO_2)""n(CO_2) = n(O_2)\/3 = 9\/3 = 3 (moleculas)"

3)

"3n(O_2) : 2n(SO_2)""n(SO_2) = (n(O_2)*2)\/3 = (9*2)\/3 = 6 (moleculas)"

4)

"n(CS_2) = n(O_2)\/3 = 9\/3 = 3 (moleculas)"

If at the beginning of the reaction the molecules CS2 was 4 , and 3 responded, respectively, after the reaction, 1 molecule CS2 remained

5) All molecules O2 reacted. After the reaction, not a single molecule O2.


Answer: 1) O2; 2) 3 moleculas; 3) 6 moleculas; 4) 3 moleculas; 5) not a single molecule.


what chemical in this balanced reaction is the limiting reactant?

Answer - O2

How many molecules of CO2 would be made on a complete reaction?

Answer - 3 molecules CO2

how many molecules of the SO2 would be made in a complete reaction?

Answer - 6 molecules SO2

how many molecules of CS2 would remain at the end of a complete reaction?

Answer - 1 molecule CS2

how many molecules of O2 would remain at the end of a complete reaction?

Answer - 0 molecules O2